Jake is 4years older than luis .In 6years ,luis age will be 2more than jakes age now.how old are they 6 years from now ?
Answer by ankor@dixie-net.com(22740)

let j = Jake's present age
let L = Luis's age
:
Write an equation for each statement
:
Jake is 4 years older than Luis .
j = L + 4
In 6 years, Luis age will be 2 more than Jake's age now.
L + 6 = j + 2
L = j + 2 - 6
L = j - 4
these two statement are the same, so we can't find their actual ages.
replace j with (L+4)
L = L + 4 - 4
L = L, not much information here
:
how old are they 6 years from now ? We don't know.